OpenStudy (anonymous):

What is the vertex of the graph of y = -4(x + 2)2 + 5? (2, 5) (-2, 5) (5, -2) (5, 2)

OpenStudy (aum):

If \(\large y = a(x-h)^2 + k\), then the vertex is: (h, k)

OpenStudy (anonymous):

ok so its A. (2,5)

OpenStudy (aum):

Compare: y = a(x-h)^2 + k to y = -4(x+2)^2 + 5 - h = +2; h = -2 +k = +5; k = 5 Vertex (-2, 5)

OpenStudy (anonymous):

oh so if the h is positive, you turn it into negative

OpenStudy (aum):

yes. k remains the same sign.

OpenStudy (anonymous):

thank you!

OpenStudy (aum):

yw.